# SlotMatch: Distilling Temporally Consistent Object-Centric Representations for Unsupervised Video Segmentation

### v2

@@ Abstract
Unsupervised video segmentation is a challenging computer vision task, especially due to the lack of supervisory signals coupled with the complexity of visual scenes. To overcome this challenge, state-of-the-art models based on slot attention often have to rely on large and computationally expensive neural architectures. To this end, we propose a simple knowledge distillation framework that effectively transfers object-centric representations to a lightweight student. The proposed framework, called SlotMatch, aligns corresponding teacher and student slots via the cosine similarity, requiring no additional distillation objectives or auxiliary supervision. The simplicity of SlotMatch is confirmed via theoretical and empirical evidence, both indicating that integrating additional losses is redundant. We conduct experiments on two datasets to compare the state-of-the-art teacher model, SlotContrast, with our distilled student. The results show that our student based on SlotMatch matches and even outperforms its teacher, while using $3.6\times$ less parameters and running $1.9\times$ faster. Moreover, our student surpasses previous unsupervised video segmentation models.
Codehttps://github.com/dianagrigore/SlotMatch
Introduction
A fundamental goal in machine perception is developing systems that, similarly to humans, understand complex visual scenes as compositions of distinct objects. This capability, studied in the area of object-centric representation learning (Burgess et al. 2019; Greff et al. 2019; Locatello et al. 2020) is a critical step for building agents that can reason about, interact with, and understand their surroundings.
Recent advances in self-supervised learning have produced powerful foundational models (Caron et al. 2021; He et al. 2022; Oquab et al. 2024). When integrated into slot-based attention frameworks (Kipf et al. 2022; Locatello et al. 2020), these models can discover and segment objects from complex scenes with remarkable fidelity. However, their success is predicated on their scale, as the high computational costs create bottlenecks for deployment. As such, typical self-supervised models are incompatible with resource-constrained environments, where object-centric representations built for video-related tasks would be most valuable.
@@ Introduction
Figure 1: Comparison of SlotContrast (teacher) versus various student versions (including our SlotMatch), showing the trade-off between performance (mBO) vs. inference speed (FPS). Circle area indicates parameter count (in millions). SlotMatch (cyan) outperforms its teacher, while being nearly twice as fast. Best viewed in color.
Our core contribution is a hybrid loss strategy that operates directly on the learned object slots. We frame the primary distillation objective as a direct matching loss, termed SlotMatch, where a slot from the student model serves as the “anchor” and the corresponding teacher slot provides the “positive” target. This pulls the student’s representation into alignment with the teacher’s proven semantic space. Crucially, the necessary repulsive force is not supplied by explicit negatives in the distillation loss itself. Instead, we rely on the student’s slot contrastive loss, which is concurrently optimized with the distillation objective. This combination effectively decomposes the learning signal: the distillation loss teaches each slot what to be, while the contrastive loss teaches it not to be a redundant copy of its other slots. We apply a similar principle to the reconstructed features, i.e. the student is trained to minimize the mean squared error between its input and output representations, without distilling representations from the teacher. We provide theoretical and empirical evidence confirming that our simple distillation objective is sufficient and effective. More precisely, we show that distilling output features is redundant.
We carry out experiments on two benchmark datasets for unsupervised video segmentation, MOVi-E (Ghorbani et al. 2021) and YTVIS-2021 (Yang et al. 2021b). We compare out student based on SlotMatch with its state-of-the-art teacher, SlotContrast (Manasyan et al. 2025), as well as other competitive methods from recent literature (Aydemir, Xie, and Guney 2023; Kipf et al. 2022; Singh, Wu, and Ahn 2022; Zadaianchuk, Seitzer, and Martius 2023). The results indicate that our student outperforms all state-of-the-art models on both datasets. As shown in Figure 1, our student contains $3.6\times$ less parameters and runs $1.9\times$ faster than its teacher model, SlotContrast.
@@
  • We empirically show that our student based on SlotMatch is both effective and efficient, surpassing its teacher in terms of multiple performance metrics, while being $1.9\times$ faster and $3.6\times$ smaller.
Related Work
Object-centric representation learning.
@@
Unsupervised scene decomposition aims to discover object-based structure in raw perceptual inputs, without supervision. Early methods approached this task via perceptual grouping (Greff et al. 2016), spatial mixture models (Greff, Van Steenkiste, and Schmidhuber 2017), or foreground-background separation (Yang et al. 2021a). Recent approaches focused on slot attention (Locatello et al. 2020), which uses iterative attention to bind latent “slots” to individual objects. Since its introduction, the paradigm has been extended to incorporate real-world images (Seitzer et al. 2023), adaptive slot counts (Fan et al. 2024), and scale-invariant pipelines (Biza et al. 2023). While all of these have shown strong results in image-based object discovery, their adaptation to video requires additional mechanisms to ensure temporal consistency.
Object-centric video models.
@@
Slot attention was originally extended to video through methods such as SAVi (Kipf et al. 2022) and SAVi++ (Elsayed et al. 2022), which introduced cross-frame attention and optical flow to promote temporal coherence. Later, STEVE (Singh, Wu, and Ahn 2022) added sequential latent dynamics for video generation. VideoSAUR (Zadaianchuk, Seitzer, and Martius 2023) and SlotContrast (Manasyan et al. 2025) further enhanced the approach with contrastive objectives and stronger backbones, e.g. DINOv2 (Oquab et al. 2024). While effective, these methods typically rely on compute-intensive encoders, limiting their applicability in real-time or embedded scenarios. Our work addresses this limitation by transferring object-centric knowledge from high-capacity video models to lightweight models via distillation.
Knowledge distillation.
@@
Knowledge distillation (KD) transfers information from a large “teacher” model to a compact “student” by aligning output distributions (Hinton, Vinyals, and Dean 2015), intermediate activations (Romero et al. 2014), or feature Jacobians (Czarnecki et al. 2017). Most of the existing KD studies are focused on traditional tasks (e.g. classification (Iordache, Alexe, and Ionescu 2025), language modeling (Gu et al. 2024)), with limited application to object-centric learning.
@@
To the best of our knowledge, there are only a few recent papers that explore distillation from scene decomposition (Kara et al. 2024; Li et al. 2024; Liao et al. 2025; Seitzer et al. 2023). DIOD (Kara et al. 2024) introduced a self-distillation strategy in video slot attention, where an EMA teacher supervises the student by refining slot masks using optical flow and static cues. DINOSAUR (Seitzer et al. 2023) is a framework that distills high-level features from a pre-trained DINOv2 encoder by training a slot-based model to reconstruct them, enabling slot emergence on real-world images. Some works (Li et al. 2024) used slot attention to distill object-centric features from an image-based detector to an event-based one via slot-aware cross-modal alignment, while others proposed federated slot learning (Liao et al. 2025), using a teacher-student decoder setup to distill shared object-centric representations across distributed domains.
@@
Method
Figure 2: Our SlotMatch framework performs knowledge distillation from a large frozen teacher model to a compact trainable student model. Both models process video frames through slot attention mechanisms, with the student learning through three loss components: reconstruction ($\mathcal{L}_{\text{rec}}$), temporal consistency ($\mathcal{L}_{\text{slot-contrast}}$), and our novel slot matching loss ($\mathcal{L}_{\text{slot-KD}}$) that directly aligns corresponding slots between teacher and student using cosine similarity. Best viewed in color.
We propose SlotMatch, a distillation framework designed to transfer temporally-consistent object-centric representations from a large teacher model to a compact student model. Our core insight is to operate directly in the slot space. As such, we introduce a novel similarity-based distillation objective that aligns semantic slot representations between teacher and student. Unlike standard methods, SlotMatch avoids conventional pixel-wise or feature-level distillation, instead relying on a direct semantic match between corresponding slots.
Problem setup.
Let $\mathbf{T}$ denote a frozen pre-trained teacher model with a high-capacity encoder, and $\mathbf{S}$ denote a lightweight trainable student. Both operate on video frames and produce a fixed number of slot representations per frame, capturing object-centric semantics. Specifically, for a video frame $\boldsymbol{x}_{t}$, the teacher and student produce the slot representations defined below:
$S^{\mathbf{T}}_{t}\!=\!\{s^{\mathbf{T}}_{t,n}\}_{n=1}^{N}\!\in\!\mathbb{R}^{N\times d},\,S^{\mathbf{S}}_{t}\!=\!\{s^{\mathbf{S}}_{t,n}\}_{n=1}^{N}\!\in\!\mathbb{R}^{N\times d},$ (1)
@@
In Figure 2, we showcase our SlotMatch framework, which operates in a dual-model fashion, with the teacher and student processing identical video inputs in parallel, through a slot-attention-based encoder-decoder architecture inspired by SlotContrast (Manasyan et al. 2025). First, a pre-trained DINOv2 (Oquab et al. 2024) backbone extracts spatial feature maps from each video frame. The teacher employs a higher-capacity encoder (e.g. ViT-B), while the student uses a lighter variant (e.g. ViT-S). Both encoders are frozen, but they contain a shallow trainable MLP that projects encoder features into a joint $d$-dimensional slot space. For each frame, we initialize a fixed number of slots $N$ and iteratively update them via attention applied to encoded features. Slot attention compresses scene information into object-centric latent vectors. In the final part, the student employs a decoder that reconstructs the original features from slots. This ensures that slots are informative and can guide unsupervised segmentation. The teacher also uses a similar decoder during its training phase, but we do not employ its decoder in the distillation stage. To avoid clutter, we refrain from depicting the decoder of the teacher in Figure 2.
@@
Training proceeds as formally described in Algorithm 1. For each mini-batch of $B$ videos with $T$ frames each, we extract per-frame slot representations $\boldsymbol{S}^{\mathbf{T}}_{b,t}$ and $\boldsymbol{S}^{\mathbf{S}}_{b,t}$ from the teacher and student models (steps 3–5), where $b$ is the video index and $t$ is the frame index. In step 7, a decoder reconstructs the input features from the student slots, producing $\hat{\boldsymbol{F}}{}^{\mathbf{S}}_{b,t}$. We next compute three loss terms: (i) a reconstruction loss $\mathcal{L}_{\text{rec}}$ that ensures the student slots retain sufficient scene-level information (step 8); (ii) a temporal contrastive loss $\mathcal{L}_{\text{slot-contrast}}$ that promotes both consistency and diversity of slots across adjacent frames (step 9); and (iii) our core distillation objective $\mathcal{L}_{\text{slot-KD}}$, which aligns each student slot with its teacher counterpart via cosine similarity (step 10). These are combined into a weighted loss (step 11) used to update the student parameters via gradient descent (step 12), while the teacher remains frozen during the whole process.
Slot-level distillation.
@@
In the teacher-student training phase, the first issue that arises is to find the correspondence between teacher and student slots, since the slots do not follow a precise order. We consider two options to determine the correspondence among teacher and student slots. One option is to perform explicit matching, e.g. via Hungarian assignment, while the other is to implicitly assume slot index correspondence, i.e. the $n$-th slot of the student corresponds to the $n$-th slot of the teacher. We ablate slot matching strategies in the supplementary and find negligible difference between matching and non-matching setups.
@@
where $B$ is the mini-batch size, $T$ is the number of frames in a video, and $N$ is the number of slots. This formulation directly distills the object-centric latent space, promoting structured and efficient knowledge transfer. Each slot from the student is serving as an anchor that must align to its corresponding teacher slot as a positive, through cosine similarity-based optimization. Our formulation creates attractive forces that guide student representations toward the teacher’s semantic space.
Temporal and reconstruction losses.
Following Manasyan et al. (2025), we integrate two auxiliary objectives to improve slot quality for the student, namely the slot-slot contrastive loss ($\mathcal{L}_{\text{slot-contrast}}$), and the feature reconstruction loss ($\mathcal{L}_{\text{rec}}$). This decomposition assigns distinct responsibilities to each loss component. Critically, the absence of explicit negative sampling in the distillation loss is compensated by the concurrent optimization of $\mathcal{L}_{\text{slot-contrast}}$, which provides the necessary repulsive forces to maintain slot distinctiveness for the student. While the slot distillation loss enforces semantic alignment between corresponding teacher-student slot pairs, the slot-slot contrastive loss maintains representational diversity among student slots, ensures temporal consistency by attracting slots representing the same object across consecutive frames, while also repelling slots from different objects within the mini-batch. The feature reconstruction loss $\mathcal{L}_{\text{rec}}$ ensures that student slots contain sufficient information to reconstruct features given by the student encoder, via its trainable decoder.
Overall training objective.
The complete training objective integrates all three components:
@@
Theoretical justification.
We conjecture that it is sufficient to employ the loss in Eq. (2) to perform effective knowledge distillation. In other words, we state that integrating additional losses, e.g. distilling the reconstructed features, is not necessary as long as $\mathcal{L}_{\text{slot-KD}}$ is minimized to zero after training. To support our conjecture, which simplifies the distillation framework, we introduce the following theorem:
@@
Method MOVi-E YTVIS-2021
Image Video Image Video
@@
SAVi (Kipf et al. 2022) - - 42.8 16.0 - - - -
STEVE (Singh, Wu, and Ahn 2022) - - 50.6 26.6 - - 15.0 19.1
VideoSAUR (Zadaianchuk, Seitzer, and Martius 2023) - - 73.9 35.6 - - 28.9 26.3
VideoSAURv2 (Manasyan et al. 2025) - - 77.1 34.4 - - 31.2 29.7
SOLV (Aydemir, Xie, and Guney 2023) 80.8 - - - - - - -
SlotContrast (Manasyan et al. 2025) (teacher) 83.9 32.4 81.7 28.6 45.5 39.7 36.8 32.4
Student (no KD) 80.1 28.9 76.6 30.1 44.5 38.8 37.2 32.1
Feature KD 81.8 31.6 80.1 28.1 45.2 38.2 35.9 31.4
Reconstruction KD 81.4 31.5 81.1 29.1 44.9 38.8 35.6 31.8
SlotMatch (MSE) 82.1 30.2 78.6 26.4 45.1 39.7 36.9 32.6
SlotMatch + Reconstruction KD 74.4 29.7 66.2 24.1 45.6 38.3 36.1 31.1
SlotMatch (predicted) 83.9 31.9 81.8 28.5 44.9 39.7 37.3 32.8
SlotMatch 84.1 33.6 81.8 30.5 45.8 39.8 36.3 32.6
Table 1: Comparison of video object segmentation performance on MOVi-E and YTVIS-2021. We report FG-ARI and mBO scores at both image and video levels. SlotMatch outperforms all baselines and ablations. The best performance on each column is highlighted in bold.
Experiments
Datasets

### v3

# SlotMatch: Distilling Object-Centric Representations for Unsupervised Video Segmentation

@@ Abstract
Unsupervised video segmentation is a challenging computer vision task, especially due to the lack of supervisory signals coupled with the complexity of visual scenes. To overcome this challenge, state-of-the-art models based on slot attention often have to rely on large and computationally expensive neural architectures. To this end, we propose a simple knowledge distillation framework that effectively transfers object-centric representations to a lightweight student. The proposed framework, called SlotMatch, aligns corresponding teacher and student slots via the cosine similarity, requiring no additional distillation objectives or auxiliary supervision. The simplicity of SlotMatch is confirmed via theoretical and empirical evidence, both indicating that integrating additional losses is redundant. We conduct experiments on three datasets to compare the state-of-the-art teacher model, SlotContrast, with our distilled student. The results show that our student based on SlotMatch matches and even outperforms its teacher, while using $3.6\times$ less parameters and running up to $2.7\times$ faster. Moreover, our student surpasses all other state-of-the-art unsupervised video segmentation models.
1 Introduction
A fundamental goal in machine perception is developing systems that, as humans, understand complex visual scenes as compositions of distinct objects. This capability, studied in the area of object-centric representation learning [Burgess-arxiv-2019, Greff-ICML-2019, Locatello-NeurIPS-2020], is a critical step for building agents that can reason about, interact with, and understand their surroundings. Recent advances in self-supervised learning have produced powerful foundational models [caron2021emerging, he2022masked, oquab2023dinov2]. When integrated into slot-based attention frameworks [kipf-ICLR-2022, Locatello-NeurIPS-2020], these models can discover and segment objects from complex scenes with remarkable fidelity. However, their success is predicated on their scale, as the high computational costs create bottlenecks for deployment. As such, typical self-supervised models are incompatible with resource-constrained environments, where object-centric representations built for video-related tasks would be most valuable.
@@ 1 Introduction
Figure 1: Comparison of SlotContrast (teacher) versus various student versions (including our SlotMatch), showing the trade-off between performance (mean Best Overlap or mBO) vs. inference speed (FPS). Circle area indicates parameter count (in millions). SlotMatch (cyan) outperforms its teacher, while being nearly twice as fast on NVIDIA A100. Best viewed in color.
Our core contribution is a novel optimization strategy that operates directly on the learned object slots. We frame the primary distillation objective as a direct matching loss, termed SlotMatch, where a slot from the student model serves as the “positive” and the corresponding teacher slot provides the fixed (target) “anchor”. This pulls the student’s representation towards the teacher’s proven semantic space. Crucially, the necessary repulsive force is not supplied by explicit negatives in the distillation loss itself. Instead, we rely on the student’s slot contrastive loss, which is concurrently optimized with the distillation objective. This combination effectively decomposes the learning signal: the distillation loss teaches each slot what to be, while the contrastive loss teaches it not to be a redundant copy of other slots. We apply a similar principle to the reconstructed features, i.e. the student is trained to minimize the mean squared error between its input and output representations, without distilling representations from the teacher. We provide theoretical and empirical evidence confirming that our simple distillation objective is sufficient and effective. More precisely, we show that distilling output features is redundant.
We carry out experiments on three benchmark datasets for unsupervised video segmentation, MOVi-E [ghorbani2021movi], YTVIS-2021 [vis2021] and DAVIS 2017 [Perazzi-CVPR-2016]. We compare out student based on SlotMatch with its state-of-the-art teacher, SlotContrast [Manasyan-CVPR-2025], as well as other competitive methods from recent literature [Aydemir-NeurIPS-2023, kipf-ICLR-2022, Singh-NeurIPS-2022, Zadaianchuk-NeurIPS-2023, Seitzer-arxiv-2022, kara2024diod]. The results indicate that our student outperforms state-of-the-art models on all datasets. Furthermore, our student contains $3.6\times$ less parameters and runs up to $2.7\times$ faster than its teacher model, SlotContrast (see Figure 1 and Table 2).
@@
  • We empirically show that our student based on SlotMatch is both effective and efficient, surpassing its teacher in terms of multiple performance metrics, while being $1.9\times$ to $2.7\times$ faster and $3.6\times$ smaller.
2 Related Work
@@
Object-centric representation learning. Unsupervised scene decomposition aims to discover object-based structure in raw perceptual inputs, without supervision. Early methods approached this task via perceptual grouping [Greff-NeurIPS-2016], spatial mixture models [Greff-NeurIPS-2017], or foreground-background separation [Yang-ICCV-2021]. Recent approaches focused on slot attention [Locatello-NeurIPS-2020], which uses iterative attention to bind latent learnable vectors, called slots, to individual objects. Since its introduction, the paradigm has been extended to incorporate real-world images [Seitzer-arxiv-2022], auto-regressive decoder with patch-order permutation [kakogeorgiou2024spot], adaptive slot counts [Fan-CVPR-2024], learned-query initialization [Jia-ICLR-2023], language-controlled slots [didolkar2025ctrl] and scale-invariant pipelines [Biza-ICML-2023]. While all of these have shown strong results in image-based object discovery, their adaptation to video requires additional mechanisms to ensure temporal consistency.
@@
Object-centric video models. Slot attention was originally extended to video through methods such as SAVi [kipf-ICLR-2022] and SAVi++ [Elsayed-NeurIPS-2022], which introduced cross-frame attention and optical flow to promote temporal coherence. Later, STEVE [Singh-NeurIPS-2022] added sequential latent dynamics for video generation. VideoSAUR [Zadaianchuk-NeurIPS-2023] and SlotContrast [Manasyan-CVPR-2025] further enhanced the approach with contrastive objectives and stronger backbones, e.g. DINOv2 [oquab2023dinov2]. While effective, these methods typically rely on computationally-intensive encoders, limiting their applicability to real-time or embedded scenarios. Our work addresses this limitation by transferring object-centric knowledge from a high-capacity video model to a lightweight model via distillation.
@@
Knowledge distillation. Knowledge distillation (KD) transfers information from a large “teacher” model to a compact “student” by aligning output distributions [Hinton-arxiv-2015], intermediate activations [Romero-arxiv-2014], or feature Jacobians [Czarnecki-NeurIPS-2017], among others. Most of the existing KD studies are focused on traditional tasks (e.g. classification [iordache2025multi], language modeling [guminillm]), with limited application to object-centric learning.
@@
To the best of our knowledge, there are only a few recent papers that explore distillation from scene decomposition [kakogeorgiou2024spot, kara2024diod, li2024object, liao2025forla, Seitzer-arxiv-2022]. DIOD [kara2024diod] introduced a self-distillation strategy in video slot attention, where an EMA teacher supervises the student by refining slot masks using optical flow and static cues. DINOSAUR [Seitzer-arxiv-2022] is a framework that distills high-level features from a pre-trained DINOv2 encoder by training a slot-based model to reconstruct them, enabling slot emergence on real-world images. Some works [li2024object] used slot attention to distill object-centric features from an image-based detector to an event-based one via slot-aware cross-modal alignment, while others proposed federated slot learning [liao2025forla], using a teacher-student decoder setup to distill shared object-centric representations across distributed domains.
@@
3 Method
Figure 2: Our SlotMatch framework performs knowledge distillation from a large frozen teacher model to a compact trainable student model. Both models process video frames through slot attention mechanisms, with the student learning through three loss components: reconstruction ($\mathcal{L}_{\text{rec}}$), temporal consistency ($\mathcal{L}_{\text{slot-contrast}}$), and our novel slot matching loss ($\mathcal{L}_{\text{slot-KD}}$) that directly aligns corresponding slots between teacher and student models using cosine similarity. Best viewed in color.
We propose SlotMatch, a distillation framework designed to transfer temporally-consistent object-centric representations from a large teacher model to a compact student model. Our core insight is to operate directly in the slot space. As such, we introduce a novel similarity-based distillation objective that aligns semantic slot representations between teacher and student models. Unlike standard methods, SlotMatch avoids conventional pixel-wise or feature-level distillation, instead relying on a direct semantic match between corresponding slots.
Problem setup. Let $\mathbf{T}$ denote a frozen pre-trained teacher model with a high-capacity encoder, and $\mathbf{S}$ denote a lightweight trainable student. Both operate on video frames and produce a fixed number of slot representations per frame, capturing object-centric semantics. Specifically, for a video frame $\boldsymbol{x}_{t}$, the teacher and the student produce the slot representations defined below:
$\boldsymbol{S}^{\mathbf{T}}_{t}\!=\!\{s^{\mathbf{T}}_{t,n}\}_{n=1}^{N}\!\in\!\mathbb{R}^{N\times d},\,\boldsymbol{S}^{\mathbf{S}}_{t}\!=\!\{s^{\mathbf{S}}_{t,n}\}_{n=1}^{N}\!\in\!\mathbb{R}^{N\times d},$ (1)
@@ 3 Method
SlotMatch framework. In Figure 2, we showcase our SlotMatch framework, which operates in a dual-model fashion, with the teacher and student models processing identical video inputs in parallel, through a slot-attention-based encoder-decoder architecture inspired by SlotContrast [Manasyan-CVPR-2025]. First, a pre-trained DINOv2 [oquab2023dinov2] backbone extracts spatial feature maps from each video frame. The teacher employs a higher-capacity encoder (e.g. ViT-B), while the student uses a lighter variant (e.g. ViT-S). Both encoders are frozen, but they contain a shallow trainable MLP that projects encoder features into a joint $d$-dimensional slot space. For each frame, we initialize a fixed number of $N$ slots and iteratively update them via attention applied to encoded features. Slot attention compresses scene information into object-centric latent vectors. In the final part, the student employs a decoder that reconstructs the original features from slots. This ensures that slots are informative and can guide unsupervised segmentation. The teacher also uses a similar decoder during its training phase, but we do not employ it in the distillation stage. To avoid clutter, we refrain from depicting the decoder of the teacher in Figure 2.
@@
Training proceeds as formally described in Algorithm 1. For each mini-batch of $B$ videos with $T$ frames each, we extract per-frame slot representations $\boldsymbol{S}^{\mathbf{T}}_{b,t}$ and $\boldsymbol{S}^{\mathbf{S}}_{b,t}$ from the teacher and student models (steps 5 and 7), where $b$ is the video index and $t$ is the frame index. In step 8, a decoder reconstructs the input features from the student slots, producing $\hat{\boldsymbol{F}}{}^{\mathbf{S}}_{b,t}$. We next compute three loss terms: (i) a reconstruction loss $\mathcal{L}_{\text{rec}}$ to ensure that student slots retain sufficient scene-level information (step 9); (ii) a temporal contrastive loss $\mathcal{L}_{\text{slot-contrast}}$ to promote both consistency and diversity of slots across adjacent frames (step 10); and (iii) our core distillation objective $\mathcal{L}_{\text{slot-KD}}$ to align each student slot with its teacher counterpart via cosine similarity (step 11). These are combined into a weighted loss (step 12) used to update the student parameters via gradient descent (step 13), while the teacher remains frozen during the whole process.
@@
Slot-level distillation. In the teacher-student training phase, the first issue that arises is to find the correspondence between teacher and student slots, since the slots do not follow a precise order. We consider two options to determine the correspondence among teacher and student slots. One option is to perform explicit matching, e.g. via Hungarian assignment, while the other is to implicitly assume slot index correspondence, i.e. the $n$-th slot of the student corresponds to the $n$-th slot of the teacher. We ablate slot correspondence strategies in Table 5 and find negligible difference between Hungarian and direct assignment setups. We thus adopt the simple direct assignment in our framework.
@@
where $B$ is the mini-batch size, $T$ is the number of frames in a video, and $N$ is the number of slots. This formulation directly distills the object-centric latent space, promoting structured and efficient knowledge transfer. Each slot from the student serves as a “positive” that must align to its corresponding teacher slot through cosine similarity-based optimization. The target teacher slot serves a fixed “anchor”, since the teacher is completely frozen during distillation. Our formulation creates attractive forces that guide the student’s representation toward the teacher’s semantic space.
Temporal and reconstruction losses. Following Manasyan-CVPR-2025, we integrate two auxiliary objectives to improve slot quality for the student, namely the slot-slot contrastive loss ($\mathcal{L}_{\text{slot-contrast}}$), and the feature reconstruction loss ($\mathcal{L}_{\text{rec}}$). This decomposition assigns distinct responsibilities to each loss component. Critically, the absence of explicit negative sampling in the distillation loss is compensated by the concurrent optimization of $\mathcal{L}_{\text{slot-contrast}}$, which provides the necessary repulsive forces to maintain slot distinctiveness for the student. While the slot distillation loss enforces semantic alignment between corresponding teacher-student slot pairs, the slot-slot contrastive loss maintains representational diversity among student slots, ensures temporal consistency by attracting slots representing the same object across consecutive frames, while also repelling slots from different objects within the mini-batch. The feature reconstruction loss $\mathcal{L}_{\text{rec}}$ ensures that student slots contain sufficient information to reconstruct features given by the student encoder, via its trainable decoder.
Overall training objective. The complete training objective integrates all three components:
@@
Theoretical justification. We conjecture that it is sufficient to employ the loss in Eq. (2) to perform effective knowledge distillation. In other words, we assume that integrating additional losses, e.g. distilling the reconstructed features, is not necessary as long as $\mathcal{L}_{\text{slot-KD}}$ is minimized to zero after training. To support our conjecture, which simplifies the distillation framework, we introduce the following theorem:
@@
Table 1: Comparison of video object segmentation performance on MOVi-E, YTVIS-2021, and DAVIS 2017. We report FG-ARI and mBO scores at both image and video levels. In most cases, SlotMatch outperforms both baselines and ablated versions. The top two scores for each metric are highlighted in blue bold (top method), orange bold (second best).
Method MOVi-E YTVIS-2021 DAVIS 2017
Image Video Image Video Image Video
@@
FG-ARI $\uparrow$
mBO $\uparrow$
FG-ARI $\uparrow$
mBO $\uparrow$
SAVi [kipf-ICLR-2022] 39.2 - 42.8 16.0 - - - - - - - -
STEVE [Singh-NeurIPS-2022] 54.1 - 50.6 26.6 - - 15.0 19.1 - - - -
VideoSAUR [Zadaianchuk-NeurIPS-2023] 78.4 - 73.9 35.6 - - 28.9 26.3 12.6 23.9 8.3 18.4
VideoSAURv2 [Manasyan-CVPR-2025] - - 77.1 34.4 - - 31.2 29.7 11.8 23.0 7.6 19.2
SOLV [Aydemir-NeurIPS-2023] 80.8 - - - 29.1 - - - 32.2 - - -
SlotContrast [Manasyan-CVPR-2025] (teacher) 83.9 32.4 81.7 28.6 45.5 39.7 36.8 32.4 89.1 12.1 73.3 11.8
DINOSAUR [Seitzer-arxiv-2022] 65.1 29.1 - - - - - - - - - -
DIOD [kara2024diod] 82.2 - - - - - - - - - - -
Student (no KD) 80.1 28.9 76.6 30.1 44.5 38.8 37.2 32.1 93.7 11.9 76.6 11.6
Feature KD 81.8 31.6 80.1 28.1 45.2 38.2 35.9 31.4 94.6 12.8 91.6 12.7
Reconstruction KD 81.4 31.5 81.1 29.1 44.9 38.8 35.6 31.8 92.0 20.5 88.4 20.3
SlotMatch (MSE) 82.1 30.2 78.6 26.4 45.1 39.7 36.9 32.6 94.4 20.0 89.6 20.0
SlotMatch + Reconstruction KD 74.4 29.7 66.2 24.1 45.6 38.3 36.1 31.1 96.4 12.7 91.2 12.4
SlotMatch (ours) 84.1 33.6 81.8 30.5 45.8 39.8 37.3 32.8 95.6 30.6 92.8 31.6
4 Experiments
4.1 Datasets